OpenStudy (anonymous):

find its domain. f(x)=x^2-3 g(x)=sqrt(21-7x) (f o g) (x)=18-7x domain of (f o g) (x) is? The domain of is

OpenStudy (turingtest):

(f o g) (x)=18-7x looks like it should be all reals, but remember that the domain of (f o g)(x) is only good when f(x) AND g(x) have that domain as well. g(x) is imaginary for x>3, and hence this limits the domain of (f o g) is \[x \le 3\]

OpenStudy (anonymous):

could you please put it in interval notation?

OpenStudy (anonymous):

nevermind i got it :) thanks so much

OpenStudy (turingtest):

\[(-\infty,3]\]

OpenStudy (anonymous):

you are pretty amazing :) could you maybe answer two more questions??? i will post them.

OpenStudy (turingtest):

try me whenever, thanks

OpenStudy (anonymous):

I'm about to post em